OpenStudy (anonymous):

Which equation shows the point-slope form of the line that passes through (3, 2) and has a slope of 1/3?

OpenStudy (er.mohd.amir):

put (3,2) at x and y which satisfy it that is eq. and find slope y=mx+c where m is slope.

OpenStudy (kidrah69):

\(\huge\color{orange}{WELCOME~TO~OPENSTUDY}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an image tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

We need to use the point slope formula \[y-y _{1}=m(x _{1}-x)\]

OpenStudy (kidrah69):

We know that m = slope our slope is 1/3 our y1= 2 and our x1=3 now plug it into the equation :)